1. Suppose you are hired by the Federal Reserve Bank of St. Louis as a financial analyst. Suppose also the tax rates and brackets are also given:

Federal Income Tax Brackets

Taxable Income

    Tax Rate

$0–9,700

10 percent

$9,701–$39,475

12 percent

$39,476–$84,200

22 percent

$84,201–$160,725

24 percent

$160,726–$204,100

$204,101–$510,300

32 percent

35 percent

Over $510,300

37 percent

  1. Calculate the total tax bill and average tax rate if your salary is $78,000
  2. Calculate the total tax bill and average tax rate if your salary is $180,000
Answer #1

ANSWER:

A) Salary is $78,000

tax upto 9,700 = tax rate * 9,700 = 10% * 9,700 = 970

tax upto (39,475 - 9,701) 29774 = tax rate * 29,774 = 12% * 29,774 = 3,572.88

tax upto (78,000 - 39,475) 38525 = tax rate * 38,525 = 22% * 38,525 = 8,475.5

total tax bill = 970 + 3,572.88 + 8,475.5 = 13,018.38

average tax rate = (total tax bill / salary) * 100 = (13,018.38 / 78,000) * 100 = 0.1669 * 100 = 16.69%

B) salary is $180,000

tax upto 9,700 = tax rate * 9,700 = 10% * 9,700 = 970

tax upto (39,475 - 9,701) 29774 = tax rate * 29,774 = 12% * 29,774 = 3,572.88

tax upto (84,200 - 39,476) 44724 = tax rate * 44,724 = 22% * 44,724 = 9,839.28

tax upto (160,725 - 84201) 76524 = tax rate * 76,524 = 24% * 76,524 = 18,365.76

tax upto (180,000 - 160,726) 19274 = tax rate * 19,274 = 32% * 19,274 = 6,167.68

total tax bill = 970 + 3,572.88 + 9,839.28 + 18,365.76 + 6,167.68 = 38,915.6

average tax rate = (total tax bill / salary) * 100 = (38,915.6 / 180,000) * 100 = 0.2161 * 100 = 21.61%